# A characterization of joint completely alternating nets

### v2

@@ Abstract.
In this article, we characterize joint completely alternating nets in terms of joint completely monotonic nets. We also discuss completely alternating sequences induced by a class of rational functions. Using the elementary techniques, we obtain a set of conditions sufficient to determine completely alternating sequences interpolated by a class of rational functions. As an application, we obtain a complete characterization of several classes of joint completely monotonic nets induced by rational functions in two variables.
@@
completely monotonic, completely alternating, Hausdorff moment2020 Mathematics Subject Classification:
@@
Let denote the set of nonnegative real numbers, set of nonnegative integers and set of natural numbers, respectively. For a net of real numbers and for let denotes the forward difference operator which is defined by
where stands for the -tuple with th entry equal to and elsewhere. When we refer as For let denote the operator and denote the sum Here, the operator refers to the identity operator i.e. Given a polynomial represents the degree of A net of real numbers is said to be a joint completely monotonic if
and joint completely alternating if
For we call the sequence as completely monotonic sequence and completely alternating sequence if it satisfies the first and second equation, respectively. Below, we enlist well-known characterizations of completely monotonic and completely alternating sequence in terms of measure (see [11, Propositions 6.11, 6.12]).
@@
Let be a positive integer and be a net of real numbers.
The net is joint completely monotonic if and only if it is a Hausdorff moment net, that is, there exists a positive Borel measure on such that
@@
The net is joint completely alternating if and only if there exist and a positive Borel measure on such that
@@
Note that a joint completely alternating net is invariant under a shift by a constant. Due to this fact, in this article, we work with joint completely alternating nets of positive real numbers.
Recall that the measure in (1.1) is called the representing measure of and it is unique. This is a consequence of the -dimensional Weierstrass theorem and the Riesz representation theorem. In [13], F. Hausdorff characterized a completely monotonic sequence by Hausdorff moment sequence and the above characterization about completely alternating sequence was first obtained as a special case of the Levy-Khinchin representation theory on abelian semigroups (see [9], [15]). For a comprehensive study on these topics, we refer the reader to [9], [11], [15], [16].
A lot of work has been done on the characterization of completely monotonic sequences in terms of Hausdorff moment sequences, for instance, [10, 12, 14] and references therein. However, the characterization of a completely monotonic sequence induced by rational functions in terms of zeroes and poles is challenging and seems to be beyond the limit of present understanding. One of the important results in this direction is the sufficient conditions to identify completely monotonic sequences introduced by Ball [7] which are solely dependent on the zeroes and poles of the rational functions. This article aims to determine a certain completely alternating sequences interpolated by rational functions with real zeros and poles in terms of sufficient conditions. In this article, we deal with the following class of rational functions, that is, for
where is a polynomial of degree at most , and ’s are distinct positive real numbers arranged in non-decreasing order.
In this article, more precisely, we concentrate on the following moment problem
@@
From the operator theoretic aspect, completely monotonic and completely alternating sequences are used in revealing the connection with subnormal and completely hyperexpansive operators, respectively (see [1], [2], [3], [4], [5], [6], [8]).
@@
Let Note that
@@
The above example motivates us to identify the completely alternating sequence among the sequence of rational functions of the form (1.3).
We now state the main results of this paper. The first one characterizes joint completely alternating nets and the second result involves a set of conditions sufficient to determine the completely alternating sequences induced by rational functions.
@@
Let be a net of positive real numbers. Then the following are equivalent
the net is joint completely alternating,
is a joint completely monotonic net.
@@
As an application of this result, we also develop a set of conditions sufficient to identify the completely alternating sequences induced by rational functions defined in (1.3), with and or , which solely depend on its zeros and poles (see Corollary 3.4, also see Theorem 3.1). These sufficient conditions are very easy to check as well as turn out to be necessary in a sense discussed in Remark 3.5. In a recent article [4], the authors have studied joint complete monotonicity of the net (see [4, Theorem 1.4]) and solved the Cauchy dual subnormality problem for a class of toral -isometric weighted -shifts which are separate -isometries. Combining Theorem 1.4 and 3.1, we also obtain a complete characterization of several classes of joint completely monotonic nets induced by rational functions in two variables (see Corollary 3.6).
@@
In this section, we present a proof of Theorem 1.4. In [5, Proposition 6] (also see [12, Theorem 2.5]), the authors have shown a correspondence between completely alternating sequences and completely monotonic sequences. It is worth mentioning here that there are only a few interesting examples of joint complete monotonic nets in more than one variable in the literature. Our result in this section allow us to construct non-trivial examples from the joint completely alternating nets (see Corollary 3.6).
For the reader’s convenience, we recall the following result which we need in the proof of Theorem 1.4 ([11, Proposition 6.10, p. 133]).
@@
Let Then the net is joint completely alternating if and only if is a joint completely monotonic net for all
@@
(i)(ii): Assume that is a joint completely alternating sequence. By Proposition 2.1, is a joint completely monotonic net, for every Equivalently, the net is joint completely monotonic for every This together with the fact that is a completely monotonic sequence yields that for every
Combining this with (2.1), we get
This shows that is a joint completely monotonic net.
(ii)(i): Assume that is a joint completely monotonic net. This together with (2.1) yields that for every
By the joint complete monotononicity of the net it is easy to see that the sequence
is completely monotonic for every Therefore its representing measure which is a weighted Lebesgue measure with weight function is positive for almost every But since is continuous on we have
This yields that for every is a joint completely monotonic net and equivalently, the net is joint completely alternating. ∎
@@
for some and We will make use of this expression throughout this section.
@@
where and Note that
In view of Theorem 1.1(ii), it suffices to check that for all for the proof of both the necessity and sufficiency part. By [4, Lemma 3.4], this is equivalent to and Further, it is equivalent to Now if then again by the partial fraction decomposition,
@@
Analogous argument, as used in the first part, works here as well.
@@
where To prove the sequence is completely alternating, by Theorem 1.1(ii), it is now sufficient to check that for This follows from Lemma 3.2. This completes the proof. ∎
@@
As an application to Theorem 1.5, we obtain the following classes of completely alternating sequences.
@@
The following corollary presents a complete characterization of certain classes of joint completely monotonic nets in two variables. This is an application of Theorems 1.4, 3.1 (also see [4][Theorems 3.1, 3.3, Question 1.1]).
@@
is a joint completely monotonic net if and only if
The net is joint completely monotonic if and only if
is a joint completely monotonic net if and only if

### v3

# A characterization of completely alternating functions

@@ Abstract.
In this article, we characterize completely alternating functions on an abelian semigroup in terms of completely monotone functions on the product semigroup . We also discuss completely alternating sequences induced by a class of rational functions and obtain a set of sufficient conditions (in terms of it’s zeros and poles) to determine them. As an application, we show a complete characterization of several classes of completely monotone functions on induced by rational functions in two variables. We also derive a set of necessary conditions for the complete monotonicity of the sequence .
@@
completely monotone, completely alternating, Hausdorff moment2020 Mathematics Subject Classification:
@@
Let be an abelian semigroup with identical involution and denotes the set of real-valued functions on For define the shift operator on by and the backward difference operator where denotes the identity operator on A function is called completely monotone if it is nonnegative and if for all finite set and
A function is called completely alternating if for all finite set and
The set of completely monotone and completely alternating functions on is denoted by and , respectively.
Given a set let denote the Cartesian product of with itself times. Let and denote the set of non-negative integers and the set of non-negative real numbers, respectively. Below, we enlist well-known characterizations of completely monotone and completely alternating functions on the abelian semigroup , in terms of measure (see [12, Propositions 6.11, 6.12, p. 134]).
@@
For the following holds:
if and only if it is a Hausdorff moment net, that is, there exists a positive Borel measure on such that
@@
if and only if there exist and a positive Borel measure on such that
@@
Note that a completely alternating function is invariant under a shift by a constant. Due to this fact, in this article, we work with completely alternating functions which are positive.
Recall that the measure in (1.1) is called the representing measure of and it is unique. This is a consequence of the -dimensional Weierstrass theorem and the Riesz representation theorem. In [14], F. Hausdorff characterized a completely monotone sequence by Hausdorff moment sequence and the above characterization about completely alternating sequence was first obtained as a special case of the Levy-Khinchin representation theory on abelian semigroups (see [10], [16]). For a comprehensive study on these topics, we refer the reader to [10], [12], [16], [17].
A lot of work has been done on the characterization of completely monotone functions in terms of Hausdorff moment sequences, for instance, [11, 13, 15] and references therein. However, the characterization of a completely monotone sequence induced by rational functions in terms of zeroes and poles is challenging and seems to be beyond the limit of present understanding. One of the important results in this direction is the sufficient conditions to identify completely monotone sequences introduced by Ball [8] which are solely dependent on the zeroes and poles of the rational functions (see Theorem 1.6). This article aims to determine certain completely alternating sequences interpolated by rational functions with real zeros and poles in terms of sufficient conditions. In this article, we deal with the following class of rational functions, that is, for
where is a polynomial of degree at most and
Here, more precisely, we concentrate on the following moment problem
@@
From the operator theoretic aspect, completely monotone and completely alternating sequences are used in revealing the connection with subnormal and completely hyperexpansive operators, respectively (see [1], [2], [3], [4], [7], [6], [9]).
@@
Let For the semigroup we write Note that
@@
The above example motivates us to identify completely alternating sequences among the sequence of rational functions of the form (1.3).
We now state the main results of this paper. The first one characterizes a completely alternating function on an abelian semigroup with the identical involution, and the second result involves a set of conditions sufficient to determine the completely alternating sequences induced by rational functions.
@@
Let S be an abelian semigroup with the identical involution and let be a strictly positive function. The function defined by
is completely monotone on the product semigroup with the identical involution if and only if is completely alternating on i.e.
@@
As an application of this result, we also develop a set of conditions sufficient to identify the completely alternating sequences induced by rational functions defined in (1.3), with and or , which solely depend on its zeros and poles (see Corollary 3.4, also see Theorem 3.1). These sufficient conditions are very easy to check. They turn out to be necessary in a sense discussed in Remark 3.5. In a recent article [4], the authors have studied complete monotonicity of the net (see [4, Theorem 1.4]) and solved the Cauchy dual subnormality problem for a class of toral -isometric weighted -shifts which are separate -isometries. Combining Theorem 1.4 and 3.1, we also obtain a complete characterization of several classes of completely monotone nets induced by rational functions in two variables (see Corollary 3.9). Also, we recover Theorem 3.1 in [4] with another characterization of completely monotone sequences as an application of Theorem 1.4 (see Corollary 3.10).
Next we turn our attention to the complete monotonicity of the following class of rational function
where and In [8], K. Ball obtain a set of sufficient conditions in terms of zeros and poles of (1.4) for which (1.4) is completely monotone. For the reader’s convenience, we recall the result here:
Theorem 1.6 (K. Ball, 1994).
To the best of our knowledge, it was not known whether the Ball’s sufficiency conditions for the complete monotonicity of (1.4) are necessary. A possible reason could be that these conditions are both necessary and sufficient for and (to be seen in Section 4). It was mentioned in his paper, that those conditions (1.5) are almost necessary (see [8, Page 3]) that is, if each is completely monotone then (1.5) hold. This follows from [8, Remark]. In Section 4, we prove that Ball’s sufficiency conditions are not necessary and obtain a new set of necessary conditions for this class which is a generatization of Ball’s sufficiency condition. Before we state this result, we need the following notation. Let denote the set of all permutations on the set .
Theorem 1.7.
Let and be two non-decreasing sequences of positive real numbers. Let and Then
If the sequence is completely monotone then there exists a with such that
(1.6) for every
If the sequence is completely alternating then there exists a with such that
for every
@@
In this section, we present a proof of Theorem 1.4. In [7, Proposition 6] (also see [13, Theorem 2.5]), the authors have shown a correspondence between completely alternating sequences and completely monotone sequences. As far as we know, there are only a few interesting examples of complete monotone functions in more than one variable in the literature. Our result in this section allows us to construct non-trivial examples from completely alternating functions (see Corollary 3.9).
For the reader’s convenience, we recall the following result which we need in the proof of Theorem 1.4 ([12, Proposition 6.10, p. 133]).
@@
Let Then if and only if for all
@@
: Assume that . By Proposition 2.1, for every Equivalently, for every This together with the fact that is completely monotone yields that for every
: Assume that This together with (2.1) yields that for every
Since
for every finite set Note that for a finite set the function This can be seen by taking in (2.2). Therefore the representing measure of the sequence is a weighted Lebesgue measure (by (2.2)) with weight function is positive for almost every But since is continuous on we have
This yields that for every and equivalently, ∎
The following result is a consequence of Theorem 1.4.
Corollary 2.2.
Assume the notations of Theorem 1.4 and Then if and only if
Proof.
This follows from Theorem 1.4 and the fact that multiplying by a positive real number does not change the completely alternating (resp. completely monotone) property of a function. ∎
@@
for some and Conversely, if where and are polynomials with and then can be written as in (3.1) (using partial fraction decomposition) with being a finite signed measure (since is finite). For irreducible with zeros in the left half-plane, one can use the method from [1, p. 800] to derive an expression like (3.1). We will make use of this expression throughout this section.
@@
where and we note that
Since is finite, the function is Lebesgue integrable on In view of Theorem 1.1(ii), it suffices to check that for all to establish both necessity and sufficiency. This condition is equivalent to requiring and Further, it is equivalent to Now if then again by the partial fraction decomposition,
@@
Analogous argument, as used in the second part, works here as well.
@@
where To prove the sequence is completely alternating, by (3.1), it is now sufficient to check that for This follows from Lemma 3.2. This completes the proof. ∎
@@
As an application of Theorem 1.5, we obtain the following classes of completely alternating sequences.
@@
The next result helps in finding more classes of completely alternating sequences where not all ’s are negative.
Proposition 3.6.
For any rational function let denote the coefficient in the partial fraction decomposition of Then
Proof.
We use induction to prove the statement. For
So the statement is true for since Assume that the statement is true for any Then
From this expression, we note that From the induction hypothesis, we obtain ∎
Now we present some classes of completely alternating sequences for which not all ’s are negative.
Corollary 3.7.
Let and with Then the sequence is completely alternating.
Proof.
Remark 3.8.
The following corollary presents a complete characterization of certain classes of completely monotone nets in two variables. This is an application of Theorems 1.4, 3.1 (also see [4][Theorems 3.1, 3.3, Question 1.1]).
@@
if and only if
if and only if
if and only if
As an immediate application of Theorem 1.4, we are able to recover Theorem 3.1 in [4] and along with that, we obtain another equivalent statement.
Corollary 3.10.
For the following are equivalent:
The net is completely monotone.
The net is completely monotone.
4. Necessary conditions
In this section, we study the necessary conditions for the complete monotonicity of (1.4). As a consequence, we also obtain necessary conditions for the sequences of the type (1.4) which are completely alternating.
We first show that Ball’s sufficiency conditions for the complete monotonicity of (1.4) are not necessary and then present a proof of Theorem 1.7 with the necessary results required to prove.
Example:
Let and For every
Let Note that
This yields that is decreasing on Since This, together with (4.1) shows that is a completely monotone sequence. Equivalently, the function is completely monotone (see [5, Proposition 6]). But this does not satisfy Ball’s sufficiency conditions.
The following lemma presents a simple characterization of the complete monotonicity of a special class. A proof in a more general setting of the following lemma has been obtained in [15, Theorem 4.1]. We present an elementary proof of the following lemma, and this will play an essential role in proving Theorem 1.7.
Lemma 4.1.
Let and Let and Then, is completely monotone if and only if
Proof.
For a proof of the sufficiency part, assume that It is easy to see that, is a sum of two completely monotone sequence, hence is completely monotone. Since the product of completely monotone sequences is completely monotone, the sequence is completely monotone. For a proof of necessity part, assume that the sequence is completely monotone. By multiplying, for any if necessary, we can assume that all ’s are distinct. Let, if possible, For every
By the partial fraction decomposition, there exists such that
Observe that Note that the representing measure of is the weighted Lebesgue measure with weight function given by
It is easy to see that there exists such that This contradicts the assumption. Hence ∎
We recall the following result which will be crucial in proving Theorem 1.7.
Proposition 4.2.
Let and be a non-decreasing sequence of positive real numbers. Let Assume that is completely monotone. Then,
Proof.
See [8, Remark]. ∎
Now, K. Ball’s sufficiency conditions combined with Propositions 4.2 and Lemma 4.1 provide a complete characterization of completely monotone rational functions of the type (1.4) in the case of
Proposition 4.3.
Let be polynomials given by and with . Then, is completely monotone if and only if
We are now ready to prove Theorem 1.7.
Proof of Theorem 1.7.
(i) We use the finite induction on . For the base cases, take and follows from Proposition 4.3. Now, assume the induction hypothesis. Let if possible there exists a such that (otherwise the result is true trivially). Clearly by Lemma 4.1. Also note that, the product of two completely monotone sequences is completely monotone. Now, by using these facts, we conclude that the sequence
is completely monotone. Now take and Similarly take and By induction hypothesis there exists such that and
Let be such that Then by (4.2), we have
By taking we have which is the desired permutation. By Proposition 4.2 and (4.3), we obtain (1.6).
(ii) This follows from (i) and the fact that if the sequence of positive real numbers is completely alternating then is completely monotone.
It might be tempting to believe that the new necessary conditions for the complete monotonicity of (1.4) are sufficient. But we now present an example which shows that the necessary conditions obtained in Theorem 1.7 are not sufficient.
Example: Let and For every
Let It is easy to verify that for all This together with (4.4) yields that is not completely monotone.